The resistance of a wire is 126.48ohms at 100°C and 100ohms at 30°C.


Determine the temperature coefficient of copper at 0°C.

1
Expert's answer
2022-02-23T10:11:01-0500

Given:

"R_1=126.48\:\Omega"

"t_1=100\:\rm {^\circ C}"

"R_2=100\:\Omega"

"t_2=30\:\rm {^\circ C}"

In general

"R(t)=R_0(1+\alpha t)""\alpha=\frac{R_1-R_2}{R_2t_1-R_1t_2}""R_0=\frac{R_1}{1+\alpha t_1}=\frac{R_1}{1+\frac{R_1-R_2}{R_2t_1-R_1t_2} t_1}"

"R_0=\frac{126.48}{1+\frac{126.48-100}{100*100-126.48*30}*100}=88.7\:\Omega"